The enthalpy of combusbon for methane at 298.15 K is -726.1 kJ mol-1 .What is its enthalpy of

formation? (given: ∆Hf (C02 ) =-393.5 kJmol-1

and ∆Hf (H2O) = -285.8 kJmol-1·.


1
Expert's answer
2022-04-27T15:01:01-0400

CH4 + 2O2 = CO2 + 2H2O ∆rH = -726.1 kJ mol-1

rH = -726.1 kJ mol-1 = ∆fH(CO2) + 2∆fH(2H2O) - 2∆fH(O2) - ∆fH(CH4) = -393.5 - 2*285.8 - 2*0 - ∆fH(CH4) = -965.1 - ∆fH(CH4)

fH(CH4) = 726.1 - 965.1 = -239 kJ mol-1